# SPECIAL RELATIVITY

### v9

@@ 1 Introduction
As known, a composition of noncollinear Lorentz boosts does not result in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the Wigner rotation (it is often called the ”Thomas rotation”). The results for the Wigner rotation in the Lorentz lab frame obtained by many experts on special relativity are incorrect. They overestimate the angle of the Wigner rotation by a relativistic factor $\gamma$ compared to its real value, and the direction of the rotation is also determined incorrectly. It should be note that the correct result was obtained in the works of several authors, which were published more than half century ago but remained unnoticed against the background of numerous incorrect works. To deduce the expression for Wigner rotation is the main purpose of Chapter 17. We demonstrate that what are usually considered the advanced parts of the theory of relativity are, in fact, quite simple.
The Wigner rotation is also relates the angular rotation velocity (Thomas precession) of the spin of an elementary particles following a curvalinear orbit. The kinematics tools needed to study the motion of charged elementary spinning particles in the storage ring are the main topic of final chapter of the book. In 1959, Bargman, Michel, and Telegdi (BMT) proposed a consistent relativistic theory for the dynamics of the spin as observed in the lab frame, which was successfully tested in experiments. It is commonly believed that the BMT equation contains the standard (and incorrect) result for the Wigner rotation in the Lorentz lab frame. There are many physicists who have already received knowledge about the Wigner (Thomas) rotation from well-known textbooks and who would say, ”The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation, of which the Wigner rotation is an integral part, and can be taken as experimental confirmations of the standard expression for the Wigner rotation” This misconception about experimental test of the incorrect expression for the Wigner rotation in the lab frame is common and pernicious. We should remark that the results in the Bargmann-Michel-Telegdi paper were obtained by the method of semi-classical approximation of the Dirac equation. The Wigner rotation was not considered at all, because the Dirac equation allow obtaining the solution for the total particle’s spin motion without an explicit splitting it into the Larmor and Wigner parts. We demonstrate that the notion that the standard (incorrect) result for the Wigner rotation as an integral part of the BMT equation in most texts is based, in turn, on an incorrect physical argument.
A large number of incorrect expressions for the Wigner (Thomas) rotation can be found in the literature. Let us discuss the reasons that underlie the derivation of the incorrect expressions. The situation relating to the use of the Wigner (Thomas) rotation theory in the relativistic spin dynamics is complex. The starting point of Bargman, Michel and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame. This brings up an interesting question: Why it is convenient to transform this equation to the rest frame as of that instant? The approach in which we deal with the proper spin is much preferred in the experimental practice due to clear physical meaning of the (3D) spin vector $\vec{s}$. Unlike (3D) momentum vector, which has definite components in each reference frame, angular momentum is defined only in one particular reference frame. It does not transform. Any statement about it refers to the rest frame as of that instant. If we say that in the lab frame the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame the spin makes this angle with the line motion of the lab frame. The analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects. In the case of spin dynamics one has no occasion to make measurements with non-inertial devices. Common textbooks presentations of the spin dynamics use the approach which deals only with observations of the inertial (lab) observer.
The conventional method used to explain the spin dynamics in the lab frame is very unusual. First, we already pointed that the laws of physics in the lab frame are able to account for all physical phenomena, including the observation made by non-inertial observer. In particular, it is possible to determine the result of spin orientation measurements with respect to the lab frame axes in the proper frame. In this case, when viewed from the lab frame, the interpretation of the spin rotation experiments in the lab frame is simple and may be derived from the prediction of the proper observer as concerns the measurement made by the lab observer. Therefore, the spin rotation measurement by a polarimeter in the lab frame is interpreted with viewpoint of the proper observer as viewing this of the lab observer. This complexity is one of the reason why authors of textbooks obtained an incorrect expression for the Wigner rotation.
It is also important to elucidate the reason why did the error in the Wigner rotation theory remain so long undetected. The main conclusion drawn from the review of the literature is as follows. Wigner (Thomas) rotation is generally presented only in the context of spin kinematics as a peculiar effect of special relativity. This is supposedly the reason why until recently no researches have noticed so serious a discrepancy between the results of different works.
We emphasize that the Wigner rotation is very basic relativistic phenomenon. It is as basic as the time dilation and length contraction. For example, we demonstrate by elementary means that relativistic (non-collinear) velocities addition is regulated by the Wigner rotation. Now let us consider the aberration theory. The theory of special relativity shows that the aberration effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of special theory of relativity. In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation. According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces. In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g. sun-based) observer have common 3-space has always been considered obvious.
@@ 2.1 What is Special Relativity?
The laws of physics are invariant with respect to Lorentz transformations. This is a restrictive principle and does not determine the exact form of the dynamics in question. Understanding the postulates of the theory of relativity is similar to understanding energy conservation: at first we learn this as a principle and later on we study microscopic interpretations that must be consistent with this principle. For any system to which the energy conservation principle can be applied, a deeper theory should exist which yields insight into the detailed physical processes involved. Of course, this deeper theory must lead to energy conservation.
The principle of conservation of energy is very useful in making analyses without knowing all the formulas of the fundamental theory. A methodological analogy with the postulates of the special relativity emerges by itself. Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame. This law would then be a phenomenological law. The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by a muon. In particular, when a Lorentz transformation of the decay law is tried, one obtains the prediction that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would be reduced to 1/2 of the origin population. We may interpret this result by saying that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$.
However, the theory of relativity is necessary incomplete. Constructive (microscopic) theories like electrodynamics or quantum field theory provide more insight into the nature of things than restrictive theories like special relativity. Relativistic kinematics is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. The point is that one can, in principle compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all. For example, muons in motion behave relativistically because the field forces that are responsible for the muon disintegration satisfy quantum field equations that are Lorentz covariant. Of course, in the ”microscopic” approach to relativistic phenomena, Lorentz covariance of all the fundamental laws of physics remains, similarly to energy conservation, an unexplained fact, but all explanation must stop somewhere.
@@ 2.2 Different Approaches to Special Relativity
Einstein formulation is based on postulates: the principle of relativity and the constancy of the velocity of light. The usual covariant formulation of the theory of relativity deals with the pseudo-Eucledian space-time geometry and with the invariance of interval $ds$, but it is understood only in a limited sense when the metric is strictly diagonal. Assuming diagonality of the metric we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
In space-time geometric approach, primary importance is attributed to the geometry of space-time; it is supposed that the geometry of space-time is a pseudo-Euclidean geometry in which only 4-tensors quantities do have real physical meaning. In this most general approach the principle of relativity in contrast to Einstein formulation of the special relativity is a simple consequence of the space-time geometry. Since the space-time geometric approach deals with all possible choices of coordinates of the chosen reference frames, the second Einstein postulate referred to the constancy of the coordinate velocity of light does not hold in this formulation of the theory of relativity. Only in Lorentz coordinates, when Einstein’s synchronization of distant clocks and Cartesian space coordinates are used, the coordinate speed of light is isotropic and constant.
@@ 2.2.1 The Usual Einstein’s Approach
The constancy of the light velocity in all inertial systems of reference is not a fundamental statement of the theory of relativity. The central principle of special relativity is the Lorentz covariance of all the fundamental laws of physics. It it important to stress at this point that the second ”postulate”, contrary to the view presented in textbooks, is not a separate physical assumption, but a convention that cannot be the subject of experimental tests.
Assuming postulate 2 on the constancy of the speed of light in all inertial frames we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations. In other words, according to such limiting understanding of the theory of relativity it is assumed that only Lorentz transformations must be used to map the coordinates of events between inertial observers.
@@ 2.2.2 The Usual Covariant Approach
In the usual covariant approach the special of relativity is understood as the theory of space-time with pseudo-Euclidean geometry. Quantities of physical interest are represented by tensors in a four-dimensional space-time, i.e. by covariant quantities, and the laws of physics are written in manifestly covariant way as four-tensor equations.
@@
To simplify our writing we will use, instead of variables $T,X,Y,Z$, variables $X^{0}=cT,~{}X^{1}=X,~{}X^{2}=Y,~{}X^{3}=Z$. Then, by adopting the tensor notation, Eq. (1) becomes $ds^{2}=\eta_{ij}dX^{i}dX^{j}$, where Einstein summation is understood. Here $\eta_{ij}$ are the Cartesian components of the metric tensor and by definition, in any Lorentz system, they are given by $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$, which is the metric canonical, diagonal form. As a consequence of the space-time geometry, Lorentz coordinates systems are connected by Lorentz transformations.
Physical quantities are represented by space-time geometric (tensor) quantities. When some basis is introduced, the representation of a tensor as geometric quantity comprise both components and basis. In usual covariant approach, one only deals with the basis components of tensors in the Lorentz coordinates i.e. with the case when the basis four-vectors are orthogonal. As a result one deals only with four-tensor equations of physics written out in the component form.
However, the concept of a tensor in the usual covariant approach is given in terms of the transformation properties of its components. For example in the usual covariant approach the electromagnetic ”tensor” $F^{\mu\nu}$ is actually not a tensor since $F^{\mu\nu}$ are only components implicitly taken in standard (orthogonal) basis. The components are coordinate quantities and they do not contain the whole information about the physical quantity, since a basis of the space-time is not included. This is no problem only in the limiting case when transformations from one orthogonal basis to another orthogonal basis are selected i.e. only assuming that Lorentz transformations must be used to map the coordinates of events. According to the usual covariant approach, another transformations from standard to non standard (not orthogonal) basis, like Galilean transformations, are ”incorrect”.
@@ 2.2.3 The Space-Time Geometric Approach
We emphasize the great freedom one has in the choice of a Minkovski space-time coordinatization. The space-time continuum, determined by the interval Eq. (1) can be described in arbitrary coordinates and not only in Lorentz coordinates. In the transition to arbitrary coordinates, the geometry of four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system. The space coordinates $x^{1},x^{2},x^{3}$ can be any quantities defining the position of particles in space, and the time coordinate $x^{0}$ can be defined by an arbitrary running clock. The components of the metric tensor in the coordinate system $x^{i}$ can be determined by performing the transformation from the Lorentz coordinates $X^{i}$ to the arbitrary variables $x^{j}$, which are fixed as $X^{i}=f^{i}(x^{j})$. One then obtains
@@
However, in the space-time geometric approach, special relativity is understood as a theory of four-dimensional space-time with pseudo-Euclidean geometry. In this formulation of the theory of relativity the space-time continuum can be described equally well from the point of view of any coordinate system, which cannot possibly change $ds$. At variance, the usual formulation of the theory of relativity also deals with the invariance of $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
Common textbook presentations of the special theory of relativity use the Einstein approach or, as generalization, the usual covariant approach which deals, as discussed above, only with components of the 4-tensors in specific (orthogonal) Lorentz basis. The fact that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is not considered in textbooks. As a consequence there is a widespread belief among experts that a transformation from an orthogonal Lorentz basis to a non orthogonal basis is incorrect, while a Lorentz transformation, which is a transformation from an orthogonal Lorentz basis to another orthogonal Lorentz basis, is correct. This is not true. We can describe physics in any arbitrary coordinates system. The different transformations of coordinates only correspond to a change in the way of components of 4-tensors are written, but not influence of 4-tensors themselves. Although the Einstein synchronization i.e. Lorentz coordinates choice, is preferred by physicists due to its simplicity and symmetry, it is nothing more ”physical” than any other. A particularly very unusual choice of coordinates, the absolute time coordinate choice, will be considered and exploited in this book.
@@ 2.3 Myth About the Incorrectness of Galilean Transformations
The use of Galilean transformations within the theory of relativity requires some special discussion. Many physicists still tend to think of Galilean transformations (which is actually a transformations from an orthogonal Lorentz basis to a non-orthogonal basis) as old, incorrect transformations between spatial coordinates and time. A widespread argument used to support the incorrectness of Galilean transformations is that they do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This idea is a part of the material in well-known books and monographs (1).
Authors of textbooks are mistaken in their belief about the incorrectness of Galilean transformations. The special theory of relativity is the theory of four-dimensional space-time with pseudo-Euclidean geometry. From this viewpoint, the principle of relativity is a simple consequence of the space-time geometry, and the space-time continuum can be described in arbitrary coordinates (2). Therefore, contrary to the view presented in most textbooks, Galilean transformations are actually compatible with the principle of relativity although, of course, they alter the form of Maxwell’s equations.
@@ 2.4 Myth About the Non-Relativistic Limit of the Lorentz Transformations
It is generally believed that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. We state that this typical textbook statement is incorrect and misleading. Kinematics is a comparative study which requires two coordinate systems, and one needs to assign time coordinates to the two systems. Different types of clock synchronization provide different time coordinates. The convention on the clock synchronization amounts to nothing more than a definite choice of the coordinate system in an inertial frame of reference in Minkowski space-time. Pragmatic arguments for choosing one coordinate system over another may therefore lead to different choices in different situations. Usually, in relativistic engineering, we have a choice between absolute time coordinate and Lorentz time coordinate. The space-time continuum can be described equally well in both coordinate systems. This means that for arbitrary particle speed, the Galilean coordinate transformations well characterize a change in the reference frame from the lab inertial observer to a comoving inertial observer in the context of the theory of relativity. Let us consider the non relativistic limit. The Lorentz transformation, for $v/c$ so small that $v^{2}/c^{2}$ is neglected can be written as $x^{\prime}=x-vt$, $t^{\prime}=t-xv/c^{2}$. This infinitesimal Lorentz transformation differs from the infinitesimal Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. The difference is in the term $xv/c^{2}$ in the Lorentz transformation for time, which is a first order term.
We only wish to emphasize here the following point. An infinitesimal Lorentz transformation differs from Galilean transformation only by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v/c$. All other higher order effects, that are Lorentz-Fitzgerald contraction, time dilation, and relativistic correction in the law of composition of velocities, can be derived mathematically, by iterating this infinitesimal transformation (3).
@@
The main difference between the Lorentz coordinatization and the absolute time coordinatization is that the transformation laws connecting coordinates and times between relatively moving systems are different. It is impossible to agree with the textbook statement that there is reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ in the non-relativistic limit. This would mean that in the non-relativistic limit infinitesimal Lorentz transformations are identical to infinitesimal Galilean transformations. This statement is absurd conclusion from a mathematical standpoint. The essence of Lorentz (or Galilean) transformations consists in their infinitesimal form: relativistic kinematics effects cannot be found by the mathematical procedure of iterating the infinitesimal Galilean transformations.

### v10

@@ 2.1 What is Special Relativity?
The laws of physics are invariant with respect to Lorentz transformations. This is a restrictive principle and does not determine the exact form of the dynamics in question. Understanding the postulates of the theory of relativity is similar to understanding energy conservation: at first, we learn this as a principle, and later on, we study microscopic interpretations that must be consistent with this principle. For any system to which the energy conservation principle can be applied, a deeper theory should exist that yields insight into the detailed physical processes involved. Of course, this deeper theory must lead to energy conservation.
The principle of conservation of energy is very useful in making analyses without knowing all the formulas of the fundamental theory. A methodological analogy with the postulates of special relativity emerges by itself. Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame. This law would then be a phenomenological law. The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by a muon. In particular, when a Lorentz transformation of the decay law is tried, one obtains the prediction that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would be reduced to 1/2 of the origin population. Here $\gamma$ is the relativistic factor and $\tau_{0}$ is the proper lifetime. We may interpret this result by saying that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$.
However, the theory of relativity is necessarily incomplete. Constructive (microscopic) theories like electrodynamics or quantum field theory provide more insight into the nature of things than restrictive theories like special relativity. Relativistic kinematics is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. The point is that one can, in principle compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all. For example, muons in motion behave relativistically because the field forces that are responsible for the muon disintegration satisfy quantum field equations that are Lorentz covariant. Of course, in the ”microscopic” approach to relativistic phenomena, Lorentz covariance of all the fundamental laws of physics remains, similarly to energy conservation, an unexplained fact, but all explanation must stop somewhere.
@@ 2.2 Different Approaches to Special Relativity
Einstein’s formulation is based on postulates: the principle of relativity and the constancy of the velocity of light. The usual covariant formulation of the theory of relativity deals with the pseudo-Euclidean space-time geometry and with the invariance of interval $ds$, but it is understood only in a limited sense when the metric is strictly diagonal. Assuming the diagonality of the metric we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
In the space-time geometric approach, primary importance is attributed to the geometry of space-time; it is supposed that the geometry of space-time is a pseudo-Euclidean geometry in which only 4-tensors quantities do have real physical meaning. In this most general approach the principle of relativity in contrast to Einstein’s formulation of special relativity is a simple consequence of space-time geometry. Since the space-time geometric approach deals with all possible choices of coordinates of the chosen reference frames, the second Einstein postulate referred to the constancy of the coordinate velocity of light does not hold in this formulation of the theory of relativity. Only in Lorentz coordinates, when Einstein’s synchronization of distant clocks and Cartesian space coordinates are used, the coordinate speed of light is isotropic and constant.
@@ 2.2.1 The Usual Einstein’s Approach
The constancy of the light velocity in all inertial systems of reference is not a fundamental statement of the theory of relativity. The central principle of special relativity is the Lorentz covariance of all the fundamental laws of physics. It is important to stress at this point that the second ”postulate”, contrary to the view presented in textbooks, is not a separate physical assumption, but a convention that cannot be the subject of experimental tests.
Assuming postulate 2 on the constancy of the speed of light in all inertial frames we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations. In other words, according to such a limiting understanding of the theory of relativity, it is assumed that only Lorentz transformations must be used to map the coordinates of events between inertial observers.
@@ 2.2.2 The Usual Covariant Approach
In the usual covariant approach, the special of relativity is understood as the theory of space-time with pseudo-Euclidean geometry. Quantities of physical interest are represented by tensors in four-dimensional space-time, i.e. by covariant quantities, and the laws of physics are written in a manifestly covariant way as four-tensor equations.
@@
To simplify our writing we will use, instead of variables $T,X,Y,Z$, variables $X^{0}=cT,~{}X^{1}=X,~{}X^{2}=Y,~{}X^{3}=Z$. Then, by adopting the tensor notation, Eq. (1) becomes $ds^{2}=\eta_{ij}dX^{i}dX^{j}$, where Einstein summation is understood. Here $\eta_{ij}$ are the Cartesian components of the metric tensor and by definition, in any Lorentz system, they are given by $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$, which is the metric canonical, diagonal form. As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations.
Physical quantities are represented by space-time geometric (tensor) quantities. When some basis is introduced, the representation of a tensor as a geometric quantity comprises both components and basis. In the usual covariant approach, one only deals with the basis components of tensors in the Lorentz coordinates i.e. with the case when the basis four-vectors are orthogonal. As a result, one deals only with four-tensor equations of physics written out in the component form.
However, the concept of a tensor in the usual covariant approach is given in terms of the transformation properties of its components. For example in the usual covariant approach the electromagnetic ”tensor” $F^{\mu\nu}$ is actually not a tensor since $F^{\mu\nu}$ are only components implicitly taken in standard (orthogonal) basis. The components are coordinate quantities and they do not contain the whole information about the physical quantity, since a basis of the space-time is not included. This is no problem only in the limiting case when transformations from one orthogonal basis to another orthogonal basis are selected i.e. only assuming that Lorentz transformations must be used to map the coordinates of events. According to the usual covariant approach, other transformations from standard to nonstandard (not orthogonal) basis, like Galilean transformations, are ”incorrect”.
@@ 2.2.3 The Space-Time Geometric Approach
We emphasize the great freedom one has in the choice of a Minkovski space-time coordinatization. The space-time continuum, determined by the interval Eq. (1) can be described in arbitrary coordinates and not only in Lorentz coordinates. In the transition to arbitrary coordinates, the geometry of four-dimensional space-time does not change, and in the special theory of relativity, we are not limited in any way in the choice of a coordinates system. The space coordinates $x^{1},x^{2},x^{3}$ can be any quantities defining the position of particles in space, and the time coordinate $x^{0}$ can be defined by an arbitrary running clock. The components of the metric tensor in the coordinate system $x^{i}$ can be determined by performing the transformation from the Lorentz coordinates $X^{i}$ to the arbitrary variables $x^{j}$, which are fixed as $X^{i}=f^{i}(x^{j})$. One then obtains
@@
However, in the space-time geometric approach, special relativity is understood as a theory of four-dimensional space-time with pseudo-Euclidean geometry. In this formulation of the theory of relativity, the space-time continuum can be described equally well from the point of view of any coordinate system, which cannot possibly change $ds$. At variance, the usual formulation of the theory of relativity also deals with the invariance of $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
Common textbook presentations of the special theory of relativity use the Einstein approach or, as generalization, the usual covariant approach which deals, as discussed above, only with components of the 4-tensors in specific (orthogonal) Lorentz basis. The fact that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is not considered in textbooks. As a consequence, there is a widespread belief among experts that a transformation from an orthogonal Lorentz basis to a nonorthogonal basis is incorrect, while a Lorentz transformation, which is a transformation from an orthogonal Lorentz basis to another orthogonal Lorentz basis, is correct. This is not true. We can describe physics in any arbitrary coordinate system. The different transformations of coordinates only correspond to a change in the way of components of 4-tensors are written, but not influence of 4-tensors themselves. Although the Einstein synchronization i.e. Lorentz coordinates choice, is preferred by physicists due to its simplicity and symmetry, it is nothing more ”physical” than any other. A particularly very unusual choice of coordinates, the absolute time coordinate choice, will be considered and exploited in this article.
@@ 2.3 Myth About the Incorrectness of Galilean Transformations
The use of Galilean transformations within the theory of relativity requires some special discussion. Many physicists still tend to think of Galilean transformations (which are transformations from an orthogonal Lorentz basis to a non-orthogonal basis) as old, incorrect transformations between spatial coordinates and time. A widespread argument used to support the incorrectness of Galilean transformations is that they do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This idea is a part of the material in well-known books and monographs 111 Many physicists tend to think of Galilean transformations as pre-relativistic transformations between spatial coordinates and time that are not compatible with the special theory of relativity. To quote e.g. Bohm [1] ”… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to the fact, the Galilean transformations evidently cannot be the correct one.”. Similar statements can also be found in recently published pedagogical papers. To quote e.g. Drake and Purvis [2] ”One of the great insights to come relativity theory was the realization that Galilean transformations are wrong. The correct way to translate the space-time measure of events between inertial frames is with the Lorentz transformations”..
Authors of textbooks are mistaken in their belief about the incorrectness of Galilean transformations. The special theory of relativity is the theory of four-dimensional space-time with pseudo-Euclidean geometry. From this viewpoint, the principle of relativity is a simple consequence of the space-time geometry, and the space-time continuum can be described in arbitrary coordinates  [^2] The mathematical argument that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is considered in textbooks as erroneous. To quote L. Landau and E. Lifshitz [3]: ”This formula is called the Galileo transformation. It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.”. This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. In pseudo-Euclidean geometry the interval between events is an invariant in arbitrary coordinates.. Therefore, contrary to the view presented in most textbooks, Galilean transformations are compatible with the principle of relativity although, of course, they alter the form of Maxwell’s equations.
@@ 2.4 Myth About the Non-Relativistic Limit of the Lorentz Transformations
It is generally believed that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. We state that this typical textbook statement is incorrect and misleading. Kinematics is a comparative study that requires two coordinate systems, and one needs to assign time coordinates to the two systems. Different types of clock synchronization provide different time coordinates. The convention on the clock synchronization amounts to nothing more than a definite choice of the coordinate system in an inertial frame of reference in Minkowski space-time. Pragmatic arguments for choosing one coordinate system over another may therefore lead to different choices in different situations. Usually, in relativistic engineering, we have a choice between absolute time coordinate and Lorentz time coordinate. The space-time continuum can be described equally well in both coordinate systems. This means that for arbitrary particle speed, the Galilean coordinate transformations well characterize a change in the reference frame from the lab inertial observer to a comoving inertial observer in the context of the theory of relativity. Let us consider the nonrelativistic limit. The Lorentz transformation, for $v/c$ so small that $v^{2}/c^{2}$ is neglected can be written as $x^{\prime}=x-vt$, $t^{\prime}=t-xv/c^{2}$. This infinitesimal Lorentz transformation differs from the infinitesimal Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. The difference is in the term $xv/c^{2}$ in the Lorentz transformation for time, which is a first-order term.
We only wish to emphasize here the following point. An infinitesimal Lorentz transformation differs from Galilean transformation only by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v/c$. All other higher order effects, that are Lorentz-Fitzgerald contraction, time dilation, and relativistic correction in the law of composition of velocities, can be derived mathematically, by iterating this infinitesimal transformation 333 There is a widespread belief that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. To quote e.g. French [4]: ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This statement is an absurd conclusion even from a mathematical standpoint. This has been recognized by some experts, perhaps most explicitly by Baierlein who states [6] that ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation. But the finite transformations are radically different, and so -however subtly-the infinitesimal Lorentz transformation must differ significantly from the Galilean transformation.” .
@@
The main difference between Lorentz coordinatization and absolute time coordinatization is that the transformation laws connecting coordinates and times between relatively moving systems are different. It is impossible to agree with the textbook statement that there is the reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ in the non-relativistic limit. This would mean that in the non-relativistic limit, infinitesimal Lorentz transformations are identical to infinitesimal Galilean transformations. The essence of Lorentz (or Galilean) transformations consists in their infinitesimal form: relativistic kinematics effects cannot be found by the mathematical procedure of iterating the infinitesimal Galilean transformations.